OpenStudy (anonymous):

Medal + Fan Please Help The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). The probability that the urine sample contains the protein albumin is 0.58 (event B). The probability that a sample contains protein, given that albumin is detected, is 1. Which statement is true? Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) =/ (B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B)=/ P(A). Events A and B are dependent because

OpenStudy (anonymous):

Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) P(B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B) P(A). Events A and B are dependent because P(A|B) = P(A).

OpenStudy (anonymous):

the spaces are suppose to have equal signs with slashes through them